# TimeRecipe: A Time-Series Forecasting Recipe via Benchmarking Module Level Effectiveness

### v1

@@ Abstract
Time-series forecasting is an essential task with wide real-world applications across domains. While recent advances in deep learning have enabled time-series forecasting models with accurate predictions, there remains considerable debate over which architectures and design components, such as series decomposition or normalization, are most effective under varying conditions. Existing benchmarks primarily evaluate models at a high level, offering limited insight into why certain designs work better. To mitigate this gap, we propose TimeRecipe, a unified benchmarking framework that systematically evaluates time-series forecasting methods at the module level. TimeRecipe conducts over 10,000 experiments to assess the effectiveness of individual components across a diverse range of datasets, forecasting horizons, and task settings. Our results reveal that exhaustive exploration of the design space can yield models that outperform existing state-of-the-art methods and uncover meaningful intuitions linking specific design choices to forecasting scenarios. Furthermore, we release a practical toolkit within TimeRecipe that recommends suitable model architectures based on these empirical insights. The benchmark is available at: https://github.com/AdityaLab/TimeRecipe.
@@ 1 Introduction
Time-series forecasting plays a critical role in a wide range of real-world domains, including economics, urban computing, and epidemiology zhu2002statstream ; zheng2014urban ; deb2017review ; mathis2024evaluation ; liu2025cape . These applications focus on predicting future trends or events based on patterns observed in historical time-series data. Recently, the emergence of deep learning has significantly advanced the field, leading to the development of numerous forecasting models lai2018modeling ; torres2021deep ; salinas2020deepar ; Yuqietal-2023-PatchTST ; zhou2021informer . These models have demonstrated strong predictive accuracy and generalization capability across diverse datasets and domains, particularly within the supervised time-series forecasting paradigm.
Despite recent successes, particularly at the model level (i.e., end-to-end forecasting architectures and pipelines), ongoing debates persist regarding the most effective deep learning strategies at the module level, referring to the internal design components within forecasting models. For example, while Transformer-based architectures are known for their ability to capture long-range temporal dependencies, they tend to struggle to generalize well on highly irregular time-series patterns, such as ETT time series. This has led to the reflection in reconsidering the effectiveness of MLP-based designs zeng2023transformers ; xu2023fits ; wang2024timemixer ; ni2025timedistill . Beyond architectural debates, further divergence arises in the design of specific modules and components. For instance, wang2024timemixer++ highlights the importance of separately modeling trend and seasonal components via series decomposition, which is overlooked in other contemporary works shi2024time . Additionally, zhao2023performative raises concerns for the effectiveness of instance normalization in short-term forecasting tasks when capturing rapid trend changes, despite it being a commonly effective practice primarily associated with long-term forecasting scenarios.
However, despite ongoing debates surrounding divergent design choices at module level, existing benchmarks only emphasize evaluation at the model level qiu2024tfb ; aksu2024gift ; li2024foundts ; du2024tsi , overlooking the importance of benchmarking the effectiveness of specific module-level components. These studies typically conclude with case-specific best-performing models, which can be less practical when applying benchmarked results to real-world forecasting scenarios that fall outside the benchmarked settings brigato2025position . Moreover, while these benchmarks are comprehensive and empirically informative, they offer limited intuitive insights into why certain models outperform others in specific forecasting scenarios, leaving important questions about model behavior and design choices underexplored.
@@
Problem Formulation. We define the time-series forecasting problem following popular existing formulations zhou2021informer ; liu2023itransformer ; liu2024time : Given historical observations $\text{X}_{t}=\{\bm{x}_{t-L},\ldots,\bm{x}_{t}\}\in\mathbb{R}^{L\times d_{\text{X}}}$ consisting of $L$ past time steps and $d_{\text{X}}$ variables at time step $t$, the goal is to predict the future $H$ steps $\text{Y}_{t}=\{\bm{x}_{t+1},\ldots,\bm{x}_{t+H}\}\in\mathbb{R}^{H\times d_{\text{X}}}$. For convenience, we denote X as the collection of all $\text{X}_{t}$ over the full time series of length $T$, and similarly, Y as the collection of all corresponding $\text{Y}_{t}$. The time-series forecasting task aims to learn a model parameterized by $\theta$ through empirical risk minimization (ERM) to obtain $f_{\theta}:\text{X}\rightarrow\text{Y}$ for all time steps $t\in T$.
@@ 2 TimeRecipe Framework
Since the introduction of Transformer-based architectures to time-series forecasting, particularly following the release of Informer zhou2021informer , most modern approaches have converged toward a common design paradigm, referred to here as the Canonical Architecture, as illustrated in Figure 1. This canonical architecture consists of five major components: pre-processing, embedding, feed-forward modeling, projection, and post-processing (which is paired with pre-processing operations). The canonical architecture serves as the foundation of TimeRecipe and captures the typical structure adopted by many state-of-the-art (SOTA) forecasting methods.
@@ 2.1 Pre-Processing
In TimeRecipe, we consider two types of popular time series pre-processing approaches that are widely used in time-series forecasting tasks: Instance Normalization kim2021reversible ; liu2022non ; fan2023dish ; liu2023adaptive ; han2024sin and Series Decomposition wu2021autoformer ; wang2024timemixer ; wang2024timemixer++ .
@@ 2.2 Embedding
In TimeRecipe, we consider four popular embedding approaches widely adopted in time-series forecasting tasks: Token zhou2021informer , Patch Yuqietal-2023-PatchTST , Invert liu2023itransformer , and Frequency xu2023fits Embedding. Additionally, we include a no-embedding variant as a controlled baseline for completeness, although it typically yields inferior performance, e.g., applying feed-forward modeling directly on the raw feature space. In short, the mathematical formulations of these embeddings are summarized in Equation 3.
$\text{X}_{t}\in[\text{B},\text{L},\text{D}],$ $\text{X}_{t}^{\text{Token}}$ $=\text{Conv}(\text{Padding}(\text{X}_{t}^{T}))^{T}$ $\in[\text{B},\text{L},\text{H}]$ (3)
@@
$\text{X}_{t}^{\text{Freq}}$ $=\text{rFFT}(\text{X}_{t})$ $\in[\text{B},\lfloor\frac{\text{L}}{2}\rfloor+1,\text{D}]$
@@
Token Embedding. Token embedding applies convolutional operations along the temporal axis of the input time series to project each timestamp into a higher-dimensional embedding space, analogous to word embeddings in natural language processing devlin2019bert . Specifically, each timestamp, comprising all features and neighboring time steps (depending on the kernel size), is treated as an individual token. This approach preserves temporal order while enabling the model to learn contextualized representations from sequences of embedded tokens.
Patch Embedding. Patch embedding segments the input time series into patches every several timestamps along the temporal dimension using convolution, and projects each patch into a higher-dimensional space, similar to the patching mechanism in vision tasks dosovitskiy2020image . In the time-series context, patch embedding operates on each feature (channel) independently. The feed-forward model then treats the collection of patches as a batch, resulting in a channel-independent processing scheme, which is a design introduced alongside the patch embedding paradigm for time-series forecasting.
@@
Frequency Embedding. Frequency embedding applies the RealFFT (rFFT) brigham2009fast to transform time-series sequences from the time domain into the frequency domain, enabling models to operate on spectral components rather than raw temporal signals. The resulting representation typically requires subsequent feed-forward models capable of handling complex-valued inputs, currently limited to MLPs in PyTorch implementations. To reconstruct the original temporal features, an inverse rFFT (irFFT) is employed in place of standard linear projections. Notably, frequency embedding is a non-parametric operation, distinguishing it from other learnable embedding methods.
@@ 2.3 Feed-Forward Modeling
Model Architectures. While Transformer-based models have achieved notable success in time-series forecasting, largely due to their ability to capture long-range dependencies across time steps or features, recent studies have also highlighted the effectiveness of MLP-based models for this task xu2023fits ; zeng2023transformers ; ni2025timedistill . Motivated by these findings, TimeRecipe aims to provide a deeper understanding of the relative effectiveness of different model architectures. For completeness, we also include RNNs as a baseline, representing a classical class of autoregressive models. RNN-based methods preceded the widespread adoption of Transformers, showing competitive performance in earlier work salinas2020deepar ; lai2018modeling and continue to demonstrate potential in more recent studies kong2025unlocking .
@@ 2.4 TimeRecipe Benchmark
IN SD Fusion Embedding FF-Type Model Venue
Feature Invert Trans. iTransformer liu2023itransformer ICLR’24
Temporal Freq. MLP FITS xu2023fits ICLR’24
Temporal Patch Trans. PatchTST Yuqietal-2023-PatchTST ICLR’23
Temporal None MLP DLinear zeng2023transformers AAAI’23
Temporal Token Trans. Informer zhou2021informer AAAI’21
Table 1: Examples of existing methods that is covered by TimeRecipe benchmark.
@@
Channel-independence Yuqietal-2023-PatchTST ; han2024capacity is an important property in time-series forecasting models, referring to the ability to forecast multiple time series independently while sharing model parameters. TimeRecipe implicitly incorporates channel-independence in its evaluations by combining specific embedding types and model architectures with temporal modeling fusion. For example, using invert embedding with an MLP model architecture alongside temporal fusion serves as an instance of channel-independent modeling.
@@
We acknowledge that TimeRecipe does not cover every existing module design in time-series forecasting approaches. The uncovered designs generally fall into two categories. First, highly specialized designs, such as tangled temporal and feature fusion with delicate design used specifically in Crossformer zhang2023crossformer . Second, designs that are of less concern in terms of module effectiveness, such as data augmentation techniques based on down-sampling wang2024timemixer ; wang2024timemixer++ , which improve the forecasting accuracy over broad experiments.
@@ 3.1 Evaluation Tasks
We distinguish different time-series forecasting tasks from two perspectives. First, from a feature perspective, we categorize tasks into multivariate forecasting, where $d_{\text{X}}>1$ and the goal is to predict the future values of all feature dimensions simultaneously, and univariate forecasting, where $d_{\text{X}}=1$ and the task focuses on forecasting future values of a single dimension based solely on its own historical observations without exogenous features. Second, from a horizon length perspective, we classify tasks into short-term and long-term forecasting. Specifically, we define short-term forecasting as cases where the lookback window is longer than the forecasting horizon (i.e., $L>H$), and long-term forecasting as cases where the lookback window is shorter than or equal to the forecasting horizon (i.e., $L\leq H$).
Following this distinction, TimeRecipe aims to comprehensively benchmark the effectiveness of the modules described in Section 2 across all types of time-series forecasting tasks, including short-term univariate, short-term multivariate, long-term univariate, and long-term multivariate forecasting problems. These tasks broadly cover a wide range of time-series forecasting scenarios, where understanding the performance of different modules under varying task settings can offer practical insights for real-world forecasting applications.
@@ 3.2 Evaluation Dataset
LTSF. The Long-Term Time-Series Forecasting (LTSF) datasets zhou2021informer are widely adopted benchmarks for evaluating both short-term and long-term forecasting tasks. They test a model’s ability to generalize across diverse domains, including Electricity Transformer Temperature (ETT; ETTh1/2, ETTm1/2) zhou2021informer , Influenza-like Illness statistics (ILI), Electricity Consumption Load (ECLasuncion2007uci , meteorological data from the National Renewable Energy Laboratory (Weather), and foreign exchange rates across various countries (Exchange).
PEMS. The Performance Measurement System (PEMS, extended Traffic dataset in LTSF) datasets li2017diffusion are standard benchmarks for time-series forecasting, commonly used in traffic prediction research. These datasets contain road occupancy or flow measurements collected by loop detectors on highways across different districts in California. We include PEMS03, PEMS04, PEMS07, and PEMS08, each varying in geographic scope, number of sensors, and data volume.
M4. The M4 dataset makridakis2018m4 is a large-scale benchmark for evaluating forecasting models across diverse real-world time series. It includes 100,000 series from domains such as macroeconomics, microeconomics, finance, industry, and demography. Each time series varies in length and frequency, spanning yearly, quarterly, monthly, weekly, daily, and hourly settings.
@@ 3.3 Evaluation Metric
3.4 Implementation and Reproducibility
Implementation. The implementation of TimeRecipe follows the widely-used Time-Series Library  [^1] https://github.com/thuml/Time-Series-Library wang2024deep . For usability, we introduce hyperparameters that control the inclusion of each module, allowing users to specify their usage of each module while TimeRecipe automatically adjusts the internal architecture and dimensions accordingly.

### v2

@@ Abstract
Time-series forecasting is an essential task with wide real-world applications across domains. While recent advances in deep learning have enabled time-series forecasting models with accurate predictions, there remains considerable debate over which architectures and design components, such as series decomposition or normalization, are most effective under varying conditions. Existing benchmarks primarily evaluate models at a high level, offering limited insight into why certain designs work better. To mitigate this gap, we propose TimeRecipe, a unified benchmarking framework that systematically evaluates time-series forecasting methods at the module level. TimeRecipe conducts over 10,000 experiments to assess the effectiveness of individual components across a diverse range of datasets, forecasting horizons, and task settings. Our results reveal that exhaustive exploration of the design space can yield models that outperform existing state-of-the-art methods and uncover meaningful intuitions linking specific design choices to forecasting scenarios. Furthermore, we release a practical toolkit within TimeRecipe that recommends suitable model architectures based on these empirical insights.
@@ 1 Introduction
Time-series forecasting plays a critical role in a wide range of real-world domains, including economics, urban computing, and epidemiology (Zhu & Shasha, 2002; Zheng et al., 2014; Deb et al., 2017; Mathis et al., 2024). These applications focus on predicting future trends or events based on patterns observed in historical time-series data. Recently, the emergence of deep learning has significantly advanced the field, leading to the development of numerous forecasting models (Lai et al., 2018; Torres et al., 2021; Salinas et al., 2020; Nie et al., 2023; Zhou et al., 2021). These models have demonstrated strong predictive accuracy and generalization capability across diverse datasets and domains, particularly within the supervised time-series forecasting paradigm.
Despite recent successes, particularly at the model level (i.e., end-to-end forecasting architectures and pipelines), ongoing debates persist regarding the most effective deep learning strategies at the module level, referring to the internal design components within forecasting models. For example, while Transformer-based architectures are known for their ability to capture long-range temporal dependencies, they tend to struggle to generalize well on highly irregular time-series patterns, such as ETT time series. This has led to the reflection in reconsidering the effectiveness of MLP-based designs (Zeng et al., 2023; Xu et al., 2023; Wang et al., 2024a; Ni et al., 2025). Beyond architectural debates, further divergence arises in the design of specific modules and components. For instance, (Wang et al., 2024a) highlights the importance of separately modeling trend and seasonal components via series decomposition, which is overlooked in other contemporary works (Shi et al., 2024). Additionally, (Zhao et al., 2025) raises concerns for the effectiveness of instance normalization in short-term forecasting tasks when capturing rapid trend changes, despite it being a commonly effective practice primarily associated with long-term forecasting scenarios.
However, despite ongoing debates surrounding divergent design choices at module level, existing benchmarks only emphasize evaluation at the model level (Qiu et al., 2024; Aksu et al., 2024; Li et al., 2024; Du et al., 2024), overlooking the importance of benchmarking the effectiveness of specific module-level components. These studies typically conclude with case-specific best-performing models, which can be less practical when applying benchmarked results to real-world forecasting scenarios that fall outside the benchmarked settings (Brigato et al., 2025). Moreover, while these benchmarks are comprehensive and empirically informative, they offer limited intuitive insights into why certain models outperform others in specific forecasting scenarios, leaving important questions about model behavior and design choices underexplored.
@@
Problem Formulation. We define the time-series forecasting problem following popular existing formulations (Zhou et al., 2021; Liu et al., 2023a; 2024a): Given historical observations $\text{X}_{t}=\{\bm{x}_{t-L},\ldots,\bm{x}_{t}\}\in\mathbb{R}^{L\times d_{\text{X}}}$ consisting of $L$ past time steps and $d_{\text{X}}$ variables at time step $t$, the goal is to predict the future $H$ steps $\text{Y}_{t}=\{\bm{x}_{t+1},\ldots,\bm{x}_{t+H}\}\in\mathbb{R}^{H\times d_{\text{X}}}$. For convenience, we denote X as the collection of all $\text{X}_{t}$ over the full time series of length $T$, and similarly, Y as the collection of all corresponding $\text{Y}_{t}$. The time-series forecasting task aims to learn a model parameterized by $\theta$ through empirical risk minimization (ERM) to obtain $f_{\theta}:\text{X}\rightarrow\text{Y}$ for all time steps $t\in T$.
@@ 2 TimeRecipe Framework
Since the introduction of Transformer-based architectures to time-series forecasting, particularly following the release of Informer (Zhou et al., 2021), most modern approaches have converged toward a common design paradigm, referred to here as the Canonical Architecture, as illustrated in Figure 1. This canonical architecture consists of five major components: pre-processing, embedding, feed-forward modeling, projection, and post-processing (which is paired with pre-processing operations). The canonical architecture serves as the foundation of TimeRecipe and captures the typical structure adopted by many state-of-the-art (SOTA) forecasting methods.
@@ 2.1 Pre-Processing
In TimeRecipe, we consider two types of popular time series pre-processing approaches that are widely used in time-series forecasting tasks: Instance Normalization (Kim et al., 2021; Liu et al., 2022c; Fan et al., 2023; Liu et al., 2023b; Han et al., 2024b) and Series Decomposition (Wu et al., 2021; Wang et al., 2024a; 2025).
@@ 2.2 Embedding
In TimeRecipe, we consider four popular embedding approaches widely adopted in time-series forecasting tasks: Token (Zhou et al., 2021), Patch (Nie et al., 2023), Invert (Liu et al., 2023a), and Frequency (Xu et al., 2023) Embedding. Additionally, we include a no-embedding variant as a controlled baseline for completeness, although it typically yields inferior performance, e.g., applying feed-forward modeling directly on the raw feature space. In short, the mathematical formulations of these embeddings are summarized in Equation 3.
$\text{X}_{t}\in[\text{B},\text{L},\text{D}],\quad$ $\text{X}_{t}^{\text{Token}}$ $=\text{Conv}(\text{Padding}(\text{X}_{t}^{T}))^{T}$ $\in[\text{B},\text{L},\text{H}]$
@@
$\text{X}_{t}^{\text{Freq}}$ $=\text{rFFT}(\text{X}_{t})$ $\in[\text{B},\lfloor\frac{\text{L}}{2}\rfloor+1,\text{D}]$ (3)
@@
Token Embedding. Token embedding applies convolutional operations along the temporal axis of the input time series to project each timestamp into a higher-dimensional embedding space, analogous to word embeddings in natural language processing (Devlin et al., 2019). Specifically, each timestamp, comprising all features and neighboring time steps (depending on the kernel size), is treated as an individual token. This approach preserves temporal order while enabling the model to learn contextualized representations from sequences of embedded tokens.
Patch Embedding. Patch embedding segments the input time series into patches every several timestamps along the temporal dimension using convolution, similar to the patching mechanism in vision tasks (Dosovitskiy et al., 2020). In the time-series context, patch embedding operates on each feature (channel) independently. The feed-forward model then treats the collection of patches as a batch, resulting in a channel-independent processing scheme, which is a design introduced alongside the patch embedding paradigm for time-series forecasting.
@@
Frequency Embedding. Frequency embedding applies the RealFFT (rFFT) (Brigham & Morrow, 2009) to transform time-series sequences from the time domain into the frequency domain, enabling models to operate on spectral components rather than raw temporal signals. The resulting representation typically requires subsequent feed-forward models capable of handling complex-valued inputs, currently limited to MLPs in PyTorch implementations. To reconstruct the original temporal features, an inverse rFFT (irFFT) is employed in place of standard linear projections. Notably, frequency embedding is a non-parametric operation, distinguishing it from other learnable embedding methods.
@@ 2.3 Feed-Forward Modeling
Model Architectures. While Transformer-based models have achieved notable success in time-series forecasting, largely due to their ability to capture long-range dependencies across time steps or features, recent studies have also highlighted the effectiveness of MLP-based models for this task (Xu et al., 2023; Zeng et al., 2023; Ni et al., 2025). Motivated by these findings, TimeRecipe aims to provide a deeper understanding of the relative effectiveness of different model architectures. For completeness, we also include RNNs as a baseline, representing a classical class of autoregressive models. RNN-based methods preceded the widespread adoption of Transformers, showing competitive performance in earlier work (Salinas et al., 2020; Lai et al., 2018) and continue to demonstrate potential in more recent studies (Kong et al., 2025).
@@ 2.4 TimeRecipe Benchmark
IN SD Fusion Embedding $F$-Type Model
Feature Invert Trans. iTransformer (Liu et al., 2023a)
Temporal Freq. MLP FITS (Xu et al., 2023)
Temporal Patch Trans. PatchTST (Nie et al., 2023), PAttn (Tan et al., 2024)
Temporal None MLP DLinear (Zeng et al., 2023)
Feature Token Trans. Autoformer (Wu et al., 2021)
Temporal Token Trans. Informer (Zhou et al., 2021)
Table 1: Examples of existing methods that are covered by TimeRecipe benchmark.
@@
Channel-independence (Nie et al., 2023; Han et al., 2024a) is an important property in time-series forecasting models, referring to the ability to forecast multiple time series independently while sharing model parameters. TimeRecipe implicitly incorporates channel-independence in its evaluations by combining specific embedding types and model architectures with temporal modeling fusion. For example, using invert embedding with an MLP model architecture alongside temporal fusion serves as an instance of channel-independent modeling.
@@
New modules can be readily incorporated into TimeRecipe. We focus on representative modules that are widely adopted and influential design choices, while omitting others to avoid the combinatorial explosion of possible configurations. The uncovered designs generally fall into two categories. First, highly specialized designs, such as tangled temporal and feature fusion with delicate design used specifically in Crossformer (Zhang & Yan, 2023). Second, designs that are of less concern on module effectiveness, such as data augmentation based on down-sampling (Wang et al., 2024a; 2025) For instance, TimeMixer (Wang et al., 2024a) can be viewed as IN+SD+Temporal-Fusion+Non-Embedding+MLP (thus falling within TimeRecipe), with an additional augmentation applied at the pre-processing stage. Augmentation approaches exhibit general benefits for predictive models with fewer concerns; therefore, they are currently omitted from our benchmark studies.
Remark 3.
While many current SOTA forecasting methods focus on foundation and prompt-based methods, TimeRecipe targets supervised learning and already includes the leading supervised approaches with existing module-level combinations. Foundation and prompt-based methods are typically built as end-to-end pre-trained frameworks, where altering a single module can disrupt the effectiveness of other pre-trained components. Furthermore, time-series foundation models remain at an early and fragmented stage, with directions ranging from scaling Transformers (e.g., Time-MOE (Shi et al., 2024), TimesFM (Das et al., )) to adapting LLMs (e.g., Chronos (Ansari et al., 2024)) or repurposing tabular foundation models (e.g., TabPFN (Hollmann et al., 2022)). Rather than benchmarking these diverse paradigms, which have already been explored in works such as GiftEval (Aksu et al., 2024), our focus is on delivering actionable insights into supervised models, which we believe can also inform the principled development of future time-series foundation models.
@@ 3.1 Evaluation Tasks
We distinguish time-series forecasting tasks from two perspectives. First, from a feature perspective, we categorize tasks into multivariate forecasting, where $d_{\text{X}}>1$ and the goal is to predict the future values of all feature dimensions simultaneously, and univariate forecasting, where $d_{\text{X}}=1$ and the task focuses on forecasting future values of a single dimension based solely on its own historical observations without exogenous features. Second, from a horizon length perspective, we classify tasks into short-term and long-term forecasting. We define short-term forecasting as cases where the lookback window is longer than the forecasting horizon (i.e., $L>H$), and long-term forecasting as cases where the lookback window is shorter than the forecasting horizon (i.e., $L\leq H$).
Following this distinction, TimeRecipe aims to comprehensively benchmark the effectiveness of the modules described in Section 2 across all time-series forecasting tasks, including short-term univariate, short-term multivariate, long-term univariate, and long-term multivariate forecasting tasks. These tasks cover a wide range of time-series forecasting scenarios, where the driven intuition on module effectivness can offer practical insights for real-world forecasting applications.
@@ 3.2 Evaluation Dataset
LTSF. The Long-Term Time-Series Forecasting (LTSF) datasets (Zhou et al., 2021) are widely adopted benchmarks for evaluating both short-term and long-term forecasting tasks. They test a model’s ability to generalize across diverse domains, including Electricity Transformer Temperature (ETT; ETTh1/2, ETTm1/2) (Zhou et al., 2021), Influenza-like Illness statistics (ILI), Electricity Consumption Load (ECL(Asuncion et al., 2007), meteorological data from the National Renewable Energy Laboratory (Weather), and foreign exchange rates across various countries (Exchange).
PEMS. The Performance Measurement System (PEMS, extended Traffic dataset in LTSF) datasets (Li et al., 2017) are standard benchmarks for time-series forecasting, commonly used in traffic prediction research. These datasets contain road occupancy or flow measurements collected by loop detectors on highways across different districts in California. We include PEMS03, PEMS04, PEMS07, and PEMS08, each varying in geographic scope, number of sensors, and data volume.
M4. The M4 dataset (Makridakis et al., 2018) is a large-scale benchmark for evaluating forecasting models across diverse real-world time series. It includes 100,000 series from domains such as macroeconomics, microeconomics, finance, industry, and demography. Each time series varies in length and frequency, spanning yearly, quarterly, monthly, weekly, daily, and hourly settings.
@@ 3.3 Evaluation Metric
Implementation. The implementation of TimeRecipe follows the widely-used Time-Series Library   [^1] https://github.com/thuml/Time-Series-Library  (Wang et al., 2024b). For usability, we introduce hyperparameters that control the inclusion of each module, allowing users to specify their usage of each module while TimeRecipe automatically adjusts the internal architecture and dimensions accordingly.